# Expert-Guided Prompting and Retrieval-Augmented Generation for Emergency Medical Service Question Answering

### v1

@@ Abstract
Large language models (LLMs) have shown promise in medical question answering, yet they often overlook the domain-specific expertise that professionals depend on—such as the clinical subject areas (e.g., trauma, airway) and the certification level (e.g., EMT, Paramedic). Existing approaches typically apply general-purpose prompting or retrieval strategies without leveraging this structured context, limiting performance in high-stakes settings. We address this gap with EMSQA, an 24.3K-question multiple-choice dataset spanning 10 clinical subject areas and 4 certification levels, accompanied by curated, subject area-aligned knowledge bases (40K documents and 2M tokens). Building on EMSQA, we introduce (i) Expert-CoT, a prompting strategy that conditions chain-of-thought (CoT) reasoning on specific clinical subject area and certification level, and (ii) ExpertRAG, a retrieval-augmented generation pipeline that grounds responses in subject area-aligned documents and real-world patient data. Experiments on 4 LLMs show that Expert-CoT improves up to 2.05% over vanilla CoT prompting. Additionally, combining Expert-CoT with ExpertRAG yields up to a 4.59% accuracy gain over standard RAG baselines. Notably, the 32B expertise-augmented LLMs pass all the computer-adaptive EMS certification simulation exams.
@@ Introduction
The rapid advancement of large language models (LLMs) has brought new possibilities to high-stakes domains such as emergency medical services (EMS) (Weerasinghe et al. 2024), where accurate and reliable decision-making is critical. There is growing interest in leveraging LLMs for medical education (Abd-Alrazaq et al. 2023), decision support (Ge et al. 2024), and certification preparation (Kung et al. 2023), particularly in the context of open-domain multiple-choice question answering (MCQA). However, while LLMs have shown promising performance on general medical QA benchmarks (Cai et al. 2024), important gaps remain between their current capabilities and the reasoning processes used by trained medical professionals.
@@
Figure 1: Overall Approach. 1) EMSQA, KB, PR Construction; 2) Tasks: Benchmark LLMs, RAG and Certification Exams.
To address these gaps, we propose a new dataset and a domain expertise-guided LLM framework that models medical MCQA in a more structured, cognitively informed manner. As shown in Figure 1, our contributions are three-fold:
@@
  • We propose two approaches to inject domain expertise into LLMs: 1) an expertise-guided prompting strategy (Expert-CoT) that encourages step-by-step reasoning from a domain-specific perspective. 2) an expertise-guided RAG method (ExpertRAG) that selectively retrieves expertise-aligned knowledge from curated EMS KBs and patient records.
@@
Medical QA datasets typically fall into two paradigms: retrieval-based tasks (Pampari et al. 2018; Krithara et al. 2023; Ben Abacha and Demner-Fushman 2019), which require explicit evidence grounding by locating answers within documents, and open-domain multiple-choice tasks such as MedQA (Jin et al. 2021), MMLU-Med (Hendrycks et al. 2021), MedMCQA (Pal, Umapathi, and Sankarasubbu 2022) test implicit medical reasoning by choosing the best answer based on world knowledge. However, existing MCQA benchmarks focus on a single certification level and provide subject area labels without corresponding knowledge bases. EMSQA is the first open-domain medical MCQA dataset that spans multiple certification levels while also furnishing both subject area annotations and a structured EMS knowledge base. Table 1 shows a comparison between EMSQA and SOTA open-domain MCQA datasets.
@@
The basic RAG framework (Lewis et al. 2020) couples a seq2seq model with a dense Wikipedia retriever, and has since been improved via query rewriting (Chan et al. 2024), entity graphs (Edge et al. 2024), and document-structure–aware retrieval (Li et al. 2025), though these methods mainly target general-domain text. For the medical domain, MedRAG (Xiong et al. 2024a) proposes a RAG pipeline with hybrid sparse–dense retrieval on MedCorps, and i-MedRAG (Xiong et al. 2024b) extends it with follow-up question generation and interactive reasoning. Self-BioRAG (Jeong et al. 2024) adapts Self-RAG (Asai et al. 2023) with domain-specific retrieval triggers, while RAG2 (Sohn et al. 2024) leverages rationale-based queries and filtering. EXPRAG (Ou et al. 2025) retrieves similar patient cases, and ClinicalRAG (Lu, Zhao, and Wang 2024) exploits medical entities to query corpora. However, existing medical RAG systems largely ignore question-specific expertise (e.g., subject area or certification level) as an explicit signal to guide retrieval and reasoning.
@@
  • Questions related to images, table were excluded. All the questions are answerable using textual inputs only.
@@
  • All questions are manually labeled with their corresponding subject areas, and both questions and answer choices have undergone human proofreading to correct any grammatical errors. A sample subset of 100 questions and KB documents were verified by an EMT expert (See Appendix Section A.5).
@@
As shown in Table 2, the dataset includes a total of 18,602 and 5,669 practice questions from public and private sources, respectively. We use the private questions exclusively for testing and split the public questions into train, validation, and test sets of 13,021, 1,860, 3,721 questions, with average token lengths of 18.27, 19.12, 18.99, respectively. More detailed statistics on preprocessed data are summarized in Appendix Section A.2.2.
@@
To build an external KB for EMSQA, we curated 16 open-access EMS education resources from reputable websites. As shown in Figure 1, these resources span four media types: Youtube video transcripts (Carrie Davis 2025; The EMS Professor 2025), official EMS guidelines (ODEMSA 2025), EMS education textbooks (American Red Cross 2025) or lecture slides (Jones & Bartlett Learning 2025), and EMS flashcards (EMT-Prep 2025). For Youtube transcripts, we segment each transcript into sections using the title cues supplied by the uploader. The PDF documents were converted to plain text with PyPDF2 and each file was parsed chapter-by-chapter using page ranges gleaned from its table of contents. We then leveraged GPT-4o (Achiam et al. 2023) to reorganize the raw chapter text into a coherent section-level hierarchy (See the prompt in Appendix Section A.3.1). Lastly, we manually audited the output by checking that every section heading and its span of text aligned with the corresponding passage in the original PDF and corrected any discrepancies. Based on the chapter name, we manually split our collected EMS KB into 10 subject areas including “airway&ventilation”, “anatomy”, “assessment”, “cardiovascular”, “ems operations”, “medical&ob”, “pediatrics”, “pharmacology”, “trauma” and “others”. After cleaning, our KB comprises 39,652 documents, contains 2,545,192 tokens with 34,110 vocabularies.
To evaluate the helpfulness of the KB for EMSQA, we assess our KB’s coverage from both syntactic and semantic aspects. For semantic evaluation, we leverage MedCPT (Jin et al. 2023) to retrieve, for each question, its most similar document from the KB and report the average similarity score between each question and its retrieved document. For syntactic evaluation, we removed stop words and computed vocabulary hit rate (KB $\cap$ QA / QA) between the KB and EMSQA. We also follow the methods in  (Ge et al. 2024) to extract EMS-related concepts defined in (Preum et al. 2020), both before and after UMLS normalization (Bodenreider 2004). Table 2 reports both syntactic and semantic overlaps with our KB. Syntactic hit rates span from 41.65% to 90.89%. Semantic similarity, measured by average cosine similarity, was 79.21% on public and 80.75% on private data. This indicates that our KB can support answering most EMSQA questions. Appendix Section A.3 provides comprehensive details on web crawling, EMS concept extraction, KB statistics, and the overlap between EMSQA and the KB.
@@
We used the National EMS Information System (NEMSIS) (Dawson 2006) 2021 public research dataset as our source of patient records (PR). NEMSIS is a large tabular corpus in which each record contains field information such as dispatch details, scene information, initial assessment, EMS protocols and triage, vital signs, medications and procedural interventions, and patient history. As shown in Figure 1, we removed any field whose value was “NA,” “Not Applicable,” “Not Recorded,” or “Unknown.” If more than 30% of a record’s fields were discarded, we excluded the entire record. Finally, we converted each remaining record into plain text by concatenating its key–value pairs. After cleaning, our corpus comprised 4,003,430 records with an average token length of 311.7. Based on the protocol field, we further categorized the patient records into 6 subject areas including “airway”, “assessment”, “cardiovascular”, “medical&ob”, “pediatrics”, and “trauma”.
As shown in Table 2, we evaluate NEMSIS coverage of EMSQA data along two dimensions: syntactic hit‐rate and semantic similarity. Semantic coverage remains high, with patient records achieving average similarity scores of 66.45% on the public split and 75.35% on the private split of EMSQA. In contrast, syntactic concept hit‐rate range from 8.87% to 28.26%, substantially lower than those observed for the KB. Detailed fields extracted from NEMSIS and their summary statistics are presented in Appendix Section A.4.
@@
Given the $i$th question $q_{i}$, its answer options $\mathcal{O}_{i}=\{o_{1},\dots,o_{m}\}$, along with the expertise annotations of subject area $s_{i}$ and certification level $l_{i}$, the goal of MCQA is to maximize the likelihood of selecting the correct answer $a_{i}^{*}\in\mathcal{O}_{i}$. Let $\mathcal{R}$ be a retriever function that takes $q_{i}$ as input and returns a set of relevant documents. A language model $f$ selects an answer $A_{i}$ as:
@@
We propose an expertise-guided LLM framework (see Figure 3) with two strategies for incorporating domain expertise in this task: (1) Expert-CoT, a prompting strategy that encodes $s_{i}$ and $l_{i}$ into a prompt template to guide LLM reasoning based on question-specific expertise; and (2) ExpertRAG, a RAG pipeline that retrieves expert-aligned knowledge conditioned on $s_{i}$. During inference, we use a trained Filter to predict subject area $\hat{s}_{i}$ and certification level $\hat{l}_{i}$ for prompt encoding and knowledge retrieval.
Filter Training
To guide the LLM reasoning and RAG retrieval based on question-specific expertise, we train a lightweight LLM-based filter to predict the key expertise attributes, including question’s subject area and certification level. As shown in Figure 3 (Left), we adopt LoRA (Hu et al. 2022) to inject a small set of trainable parameters into the model while keeping the full LLM weights fixed. We augment the LoRA modules with two classification heads, $W_{\mathrm{sub}}$ and $W_{\mathrm{lvl}}$, that predict the question’s subject area and certification level, and optimize them jointly in a multi‐task setting. We append a special token <classify> at the end of each query, and extract the hidden state $h_{i}$ of this final token from the last layer and feed it into our two classification heads:
@@
where $\sigma$ is sigmoid function. We use binary cross‐entropy for subject area classification and cross‐entropy for certification classification. We set hyper-parameters $w_{\mathrm{sub}}$ and $w_{\mathrm{lvl}}$ to balance the multi-task loss. The overall training objective is
@@
RQ3: Can expertise-aware LLMs meet the NREMT passing score at different certification levels?
@@
We use three categories of SOTA baseline models: 1) Open-source LLMs, we select Qwen3-32B (Team 2025), and LLama-3.3-70B (Grattafiori et al. 2024) because of their great performance in multiple domains; 2) Medical LLMs, we select OpenBioLLM-70B (Ankit Pal 2024), which currently leads the Open Medical-LLM Leaderboard (Pal et al. 2024). 3) Closed-source LLMs, we choose OpenAI-o3 (Brown et al. 2020) and Gemini-2.5-pro (Team et al. 2023), both of which achieve top results on a range of benchmarks. We further apply 0 to 64-shot and CoT prompting  (Wei et al. 2022) and Expert-CoT to each baseline to benchmark the performance under different prompt strategies.
@@
We select the following SOTA medical RAG models due to their superior performance and code availability: 1) MedRAG (Xiong et al. 2024a), which is a RAG toolkit combining multiple medical documents; 2) i-MedRAG (Xiong et al. 2024b), which iteratively refines medical queries via multi-step retrieval; 3) Self-BioRAG (Jeong et al. 2024), which self-reflectively decides when to retrieve biomedical texts and then generates the answer; 4) Qwen3-4B + KB, which is a vanilla RAG pipeline with our collected KB as retrieval corpora; 5) Qwen3-4B + PR, a vanilla RAG pipeline with our collected PR as retrieval corpora; 6) Qwen3-4B + Global, a vanilla RAG pipeline with both KB and PR as retrieval corpora. We also include 7) Qwen3-4B with 0-shot and 8) Qwen3-4B with CoT prompting as baselines. For a fair comparison, we applied RAG with Qwen3-4B across all methods, except Self-BioRAG, which is trained from scratch. All baseline RAG methods used CoT prompting.
@@
For ExpertRAG, we use Qwen3-4B as the core LLM. We employ MedCPT (Jin et al. 2023) as our retriever due to its superior performance in the medical domain and its widespread use in SOTA RAG systems. We fix the number of retrieved documents to $M=32$ for KB retrieval and $N=8$ for PR retrieval. We chunk all KB and PR documents using a fixed window size of 512 tokens with an overlap of 128 tokens. Since some questions in our MCQA dataset have multiple correct answers, we report both exact-match accuracy (Acc) and sample-based F1 over all answer options (F1) (Khashabi et al. 2018; Wang et al. 2019).
To train the filter, we fine-tune LoRA modules with rank r=8r{=}8, scaling factor $\alpha{=}16$, and a dropout rate of 0.05, using the sequence length of 128 tokens. To balance the certification and subject area classification objectives, we apply DWA (Liu, Johns, and Davison 2019) with $T=2$ to dynamically adjust $w_{\mathrm{cat}}$ and $w_{\mathrm{lvl}}$. We use AdamW (Loshchilov and Hutter 2017) optimizer and regularization with a weight decay of 0.01. We set the decision threshold of 0.5 for subject area classification. We fix the random seed at 42, and run all experiments on NVIDIA H200 GPUs.
Experimental Results
Filter Performance
Since subject area classification is a multi-label task and certification level classification is a multi-class task, we evaluate the classification by micro f1-score (miF) and macro f1-score (maF). The performance of filter and other baselines (0-shot, 4-shot, and CoT) are shown at Table 3. Results show our filter trained with LoRA with two classification heads significantly outperforms the baselines.
Split Model Method Subject Area Certification
miF maF miF maF
Public Filter LoRA 80.72 71.92 65.87 63.45
Qwen3-4B 0-shot 55.43 51.61 45.77 30.49
Qwen3-4B 4-shot 56.33 54.42 45.46 30.28
Qwen3-4B CoT 59.72 55.66 47.80 35.77
Private Filter LoRA 79.06 70.48 65.54 63.50
Qwen3-4B 0-shot 42.93 31.73 44.12 25.01
Qwen3-4B 4-shot 45.76 34.08 44.04 29.41
Qwen3-4B CoT 46.22 35.49 47.70 31.92
Table 3: Filter’s Performance on Expertise Classification
LLM Benchmarking and Expert-CoT Evaluation
To evaluate the strengths and limitations of SOTA LLMs on EMSQA (RQ1) and to assess how domain expertise, injected via Expert-CoT, influences reasoning (RQ2), we benchmark multiple LLMs under different prompting strategies. Table 4 and Figure 4 present the results. We highlight several key findings from this evaluation:
@@
Table 5: End-to-end RAG Performance and Ablation Study on ExpertRAG and Expert-CoT: “GT” = Ground-truth expertise; “Filter” = Predicted expertise.
@@
Few-shot prompting improves accuracy up to a point. We varied the number of in‑context exemplars from 0 to 64 (See Appendix Section A.7.2) and observed that incorporating a small number of examples yields substantial gain over the zero‑shot baseline. However, adding examples beyond a certain point leads to diminishing or no improvement.
Expert-CoT help guide LLM reasoning. Integrating domain expert knowledge via CoT‑Expert guides reasoning towards appropriate context and consistently boosts CoT prompting performance by up to 2.05% across models. Also, using the Filter’s predicted expertise for Expert-CoT leads to comparable performance to when using ground-truth.
@@
Baseline LLMs underperform on easier questions. As shown in Figure 4a, across all models, we observe that performance is lowest for the EMR certification level (the most basic tier in the NREMT exam) and highest for the Paramedic level. This may be due to smaller data size for EMR level and the procedural nature of EMR questions, whereas Paramedic questions aligning more closely with the academic-style content seen during LLM pretraining.
@@
LLMs excel in pharmacology and anatomy but falter on core NREMT domains. Figure 4b shows that models reliably answer pharmacology and anatomy questions but perform poorly on trauma, operations, and medical&OB. One reason may be task complexity. The questions in the former subject areas need single-hop, fact-based queries which can be easily solved in a zero-shot manner, whereas the latter demand multi-hop reasoning and richer EMS context.
Ablation Study on ExpertRAG and Expert-CoT
@@
To study how much does expertise help LLMs (RQ2), we report an ablation over six configurations at Table 5: (1) No-RAG, (2) RAG+CoT with a standard global retriever on EMS PR and KB, (3) RAG+Expert-CoT, (4) ExpertRAG-GT+CoT, (5) ExpertRAG-GT+Expert-CoT, and (6) ExpertRAG-Filter+Expert-CoT. The best configuration outperforms the baseline by 4.59 / 3.69 points in Acc / F1.
@@
Effect of PR and KB. Row 2 vs. row 1 isolates the gain of adding PR and KB as retrieval documents with standard global retriever. Results show KB brings more improvement than PR, and combing both yields the best performance.
@@
Effect of Expert-CoT. Row 3 vs. row 2 ablates the additional gain from Expert-CoT, showing that expertise-aware reasoning is better than standard reasoning.
@@
Effect of Expert-RAG. Rows 4 vs. 2 (and 5 vs. 3) ablate the impact of our Expert-RAG (FTR/RTF) compared with standard RAG with a global retriever. With ground-truth subject area and certification level, ExpertRAG consistently outperforms SOTA RAG baselines, highlighting the value of expertise-guided retrievers. Across settings, both FTR and RTF outperform global retrieval, and RTF achieves the best overall performance.

### v2

@@ Abstract
Large language models (LLMs) have shown promise in medical question answering, yet they often overlook the domain-specific expertise that professionals depend on-such as the clinical subject areas (e.g., trauma, airway) and the certification level (e.g., EMT, Paramedic). Existing approaches typically apply general-purpose prompting or retrieval strategies without leveraging this structured context, limiting performance in high-stakes settings. We address this gap with EMSQA, an 24.3K-question multiple-choice dataset spanning 10 clinical subject areas and 4 certification levels, accompanied by curated, subject area-aligned knowledge bases (40K documents and 2M tokens). Building on EMSQA, we introduce (i) Expert-CoT, a prompting strategy that conditions chain-of-thought (CoT) reasoning on specific clinical subject area and certification level, and (ii) ExpertRAG, a retrieval-augmented generation pipeline that grounds responses in subject area-aligned documents and real-world patient data. Experiments on 4 LLMs show that Expert-CoT improves up to 2.05% over vanilla CoT prompting. Additionally, combining Expert-CoT with ExpertRAG yields up to a 4.59% accuracy gain over standard RAG baselines. Notably, the 32B expertise-augmented LLMs pass all the computer-adaptive EMS certification simulation exams.
@@ Introduction
The rapid advancement of large language models (LLMs) has brought new possibilities to high-stakes domains such as emergency medical services (EMS) (Weerasinghe et al. 2024), where accurate and reliable decision-making is critical. There is growing interest in leveraging LLMs for medical education (Abd-Alrazaq et al. 2023), decision support (Ge et al. 2024; Luo et al. 2025), and certification preparation (Kung et al. 2023), particularly in the context of open-domain multiple-choice question answering (MCQA). However, while LLMs have shown promising performance on general medical QA benchmarks (Cai et al. 2024), important gaps remain between their current capabilities and the reasoning processes used by trained medical professionals.
@@
Figure 1: Overall Approach. (1) EMSQA, KB, PR Construction; (2) Tasks: Benchmark LLMs, RAG and Certification Exams.
To address these gaps, we propose a new dataset and a domain-expertise-guided LLM framework that models medical MCQA in a more structured, cognitively informed manner. As shown in Figure 1, our contributions are three-fold:
@@
  • We propose a expertise-guided LLM framework that infers the domain expertise attributes and injects them into LLMs using two approaches: (1) an expertise-guided prompting strategy (Expert-CoT) that encourages step-by-step reasoning from a domain-specific perspective. (2) an expertise-guided RAG method (ExpertRAG) that selectively retrieves expertise-aligned knowledge from curated EMS KBs and patient records.
@@
Medical QA datasets typically fall into two paradigms: retrieval-based tasks (Pampari et al. 2018; Krithara et al. 2023; Ben Abacha and Demner-Fushman 2019), which require explicit evidence grounding by locating answers within documents, and open-domain multiple-choice tasks such as MedQA (Jin et al. 2021), MMLU-Med (Hendrycks et al. 2021), and MedMCQA (Pal, Umapathi, and Sankarasubbu 2022) that test implicit medical reasoning by choosing the best answer based on world knowledge. However, existing MCQA datasets focus on a single certification level and provide subject area labels without corresponding knowledge bases. EMSQA is the first open-domain medical MCQA dataset that spans multiple certification levels while also furnishing both subject area annotations and a structured EMS knowledge base. Table 1 shows a comparison between EMSQA and SOTA open-domain MCQA datasets.
@@
The basic RAG framework (Lewis et al. 2020) couples a seq2seq model with a dense Wikipedia retriever, and has since been improved via query rewriting (Chan et al. 2024), entity graphs (Edge et al. 2024), and document-structure–aware retrieval (Li et al. 2025a), though these methods mainly target general-domain text. For the medical domain, MedRAG (Xiong et al. 2024a) proposes a RAG pipeline with hybrid sparse–dense retrieval on MedCorps, and i-MedRAG (Xiong et al. 2024b) extends it with follow-up question generation and interactive reasoning. Self-BioRAG (Jeong et al. 2024) adapts Self-RAG (Asai et al. 2023) with domain-specific retrieval triggers, while RAG2 (Sohn et al. 2024) leverages rationale-based queries and filtering. EXPRAG (Ou et al. 2025) retrieves similar patient cases, and ClinicalRAG (Lu, Zhao, and Wang 2024) exploits medical entities to query corpora. However, existing medical RAG systems largely ignore question-specific expertise (e.g., subject area or certification) as explicit signals to guide retrieval and reasoning. Unlike Metadata-RAG (Oudenhove 2024), which parses metadata directly from the query, we inject expertise attributes inferred from the question by a model trained on our Q&A dataset.
@@
  • Questions related to images and tables were excluded. All the questions are answerable using text inputs only.
@@
  • All questions are manually labeled with subject areas, and both questions and answer choices have undergone human proofreading to correct any grammatical errors. A sample of 100 questions and KB documents was further verified by an EMT expert (see Appendix A.5).
@@
As shown in Table 2, the dataset includes a total of 18,602 and 5,669 practice questions from public and private sources, respectively. We use the private questions exclusively for testing and split the public questions into train, validation, and test sets of 13,021, 1,860, 3,721 questions, with average token lengths of 18.27, 19.12, 18.99, respectively. More detailed statistics are provided in Appendix A.2.2.
@@
To build an external KB for EMSQA, we curated 16 open-access EMS education resources from reputable websites. As shown in Figure 1, these resources span four media types: YouTube video transcripts (Carrie Davis 2025; The EMS Professor 2025), official EMS guidelines (ODEMSA 2025), EMS education textbooks (American Red Cross 2025) or lecture slides (Jones & Bartlett Learning 2025), and EMS flashcards (EMT-Prep 2025). We segmented each YouTube transcript into sections using the title cues supplied by the uploader. The PDF documents were converted to plain text with PyPDF2 (Fenniak et al. 2022) and split to chapters using page ranges from their tables of contents. We then leveraged GPT-4o (Achiam et al. 2023) to reorganize the raw chapter text into a coherent section-level hierarchy (See Appendix A.3.1). Finally, we manually audited each section to ensure its heading and text span aligned with the corresponding passage in the original PDF and corrected any discrepancies. Due to the lack of certification information in the sources, we only categorized the chapters into 10 subject areas based on their titles: “airway&ventilation”, “anatomy”, “assessment”, “cardiovascular”, “ems operations”, “medical&ob”, “pediatrics”, “pharmacology”, “trauma” and “others”. Our final EMS KB comprises 39,652 sections, totaling 2,545,192 tokens and 34,110 unique vocabularies.
To evaluate the helpfulness of the KB for EMSQA, we assess our KB’s coverage from both syntactic and semantic aspects. For semantic evaluation, we leverage MedCPT (Jin et al. 2023) to retrieve, for each question, its most similar document from the KB and report the average similarity score between each question and its retrieved document. For syntactic evaluation, we removed stop words and computed vocabulary hit rate (KB $\cap$ QA / QA) between the KB and EMSQA. We also follow the methods in  (Ge et al. 2024) to extract EMS-related concepts defined in (Preum et al. 2020), both before and after UMLS normalization (Bodenreider 2004). Table 2 reports both syntactic and semantic overlaps with our KB. Syntactic hit rates span from 41.65% to 90.89%. Semantic similarity, measured by average cosine similarity, was 79.21% on public and 80.75% on private data. This indicates that our KB can support answering most EMSQA questions. Appendix A.3 provides comprehensive details on web crawling, EMS concept extraction, KB statistics, and the overlap between EMSQA and the KB.
@@
We used the National EMS Information System (NEMSIS) (Dawson 2006) 2021 public research dataset as our source of patient records (PR). NEMSIS is a large tabular corpus in which each record includes fields such as dispatch details, scene information, initial assessment, EMS protocols and triage, vital signs, medications and procedural interventions, and patient history. As shown in Figure 1, we removed any field whose value was “NA,” “Not Applicable,” “Not Recorded,” or “Unknown.” If more than 30% of a record’s fields were discarded, we excluded the entire record. Finally, we converted each remaining record into plain text by concatenating its key–value pairs and categorized the patient records into 6 subject areas based on their protocol field: “airway”, “assessment”, “cardiovascular”, “medical&ob”, “pediatrics”, and “trauma”. Our final corpus comprises 4,003,430 records with an average token length of 311.7. As shown in Table 2, semantic coverage of NEMSIS over EMSQA is high, with patient records reaching 66.45%/75.35% similarity on the public/private split. In contrast, syntactic concept hit rates are much lower (8.87%–28.26%) than those of the KB. Extracted NEMSIS fields and summary statistics are provided in Appendix A.4.
@@
Given the $i$th question $q_{i}$ and its answer options $\mathcal{O}_{i}=\{o_{1},\dots,o_{m}\}$, the goal of MCQA is to maximize the likelihood of selecting the correct answer $a_{i}^{*}\in\mathcal{O}_{i}$. Let $\mathcal{R}$ be a retriever that takes $q_{i}$ as input and returns a set of relevant documents. A language model $f$ selects an answer $a_{i}$ as:
@@
We propose an expertise-guided LLM framework (see Figure 3) with an expertise classification module (called Filter), which infers the domain expertise attributes of subject area $s_{i}$ and certification level $l_{i}$ from the input question $q_{i}$, and incorporates them into $f$ using two strategies: (1) Expert-CoT, a prompting strategy that encodes $s_{i}$ and $l_{i}$ into a prompt template to guide $f$ based on question-specific expertise; and (2) ExpertRAG, a subject-area-specific $\mathcal{R}$ that retrieves knowledge sources conditioned on $s_{i}$.
Filter Design
To guide the LLM reasoning and RAG retrieval based on question-specific expertise, we train a lightweight LLM-based filter to infer the key expertise attributes, including question’s subject area and certification level. As shown in Figure 3 (Left), we adopt LoRA (Hu et al. 2022) to inject a small set of trainable parameters into the model while keeping the full LLM weights fixed. We augment the LoRA modules with two classification heads, $W_{\mathrm{sub}}$ and $W_{\mathrm{lvl}}$, that predict the question’s subject area and certification level, and optimize them jointly in a multi‐task setting (Li et al. 2025b). We append a special token <classify> at the end of each query, and extract the hidden state $h_{i}$ of this final token from the last layer and feed it into our two classification heads:
@@
where $\sigma$ is sigmoid function. We use binary cross‐entropy for subject area classification and cross‐entropy for certification classification. We set hyper-parameters $w_{\mathrm{sub}}$ and $w_{\mathrm{lvl}}$ to balance the multi-task loss. The overall training loss is:
@@
RQ3: Can expertise-aware LLMs pass the NREMT standardized tests at different certification levels?
@@
We use three categories of SOTA baseline models: (1) Open-source LLMs: we select Qwen3-32B (Team 2025), and LLama-3.3-70B (Grattafiori et al. 2024) because of their great performance in multiple domains; (2) Medical LLMs: we select OpenBioLLM-70B (Ankit Pal 2024), which currently leads the Open Medical-LLM Leaderboard (Pal et al. 2024). (3) Closed-source LLMs: we choose OpenAI-o3 (Brown et al. 2020) and Gemini-2.5-pro (Team et al. 2023), both of which achieve top results on a range of benchmarks. We further apply 0- to 64-shot, CoT (Wei et al. 2022), and Expert-CoT prompting to each baseline to benchmark the performance across prompt strategies.
@@
We select the following SOTA medical RAG models due to their superior performance and code availability: (1) MedRAG (Xiong et al. 2024a), which is a RAG toolkit combining multiple medical documents; (2) i-MedRAG (Xiong et al. 2024b), which iteratively refines medical queries via multi-step retrieval; (3) Self-BioRAG (Jeong et al. 2024), which self-reflectively decides when to retrieve biomedical texts and then generates the answer; (4) Qwen3-4B + KB, which is a vanilla RAG pipeline with our collected KB as retrieval corpora; (5) Qwen3-4B + PR, a vanilla RAG pipeline with our collected PR as retrieval corpora; (6) Qwen3-4B + Global, a vanilla RAG pipeline with both KB and PR as retrieval corpora. We also include (7) Qwen3-4B with 0-shot and (8) Qwen3-4B with CoT prompting as baselines. For a fair comparison, we applied RAG with Qwen3-4B across all methods, except Self-BioRAG, which is trained from scratch. All baseline RAG methods used CoT prompting.
@@
For Expert-RAG, we use Qwen3 as the core LLM, as it is the best-performing open-source model in our benchmarking. We employ MedCPT as our retriever due to its strong performance in medical domain and widespread use in SOTA RAG. We fix the number of retrieved documents at $M=32$ for KB retrieval and $N=8$ for PR retrieval. All KB and PR documents are chunked with a window of 512 tokens with an overlap of 128 tokens. Since some questions in EMSQA have multiple correct answers, we report both exact-match accuracy (Acc) and sample-based F1 (Khashabi et al. 2018).
ßTo train the filter, we fine-tune LoRA modules with rank r=8r{=}8, scaling factor $\alpha{=}16$, and a dropout rate of 0.05, using the sequence length of 128 tokens. To balance the certification and subject area classification objectives, we apply DWA (Liu, Johns, and Davison 2019) with $T=2$ to dynamically adjust $w_{\mathrm{cat}}$ and $w_{\mathrm{lvl}}$. We use AdamW (Loshchilov and Hutter 2017) optimizer and regularization with a weight decay of 0.01. We set the decision threshold of 0.5 for subject area classification. We fix the random seed at 42, and run all experiments on NVIDIA H200 GPUs.
@@
Experimental Results
LLM Benchmarking
To evaluate the overall performance of SOTA LLMs on EMSQA (RQ1), we benchmark multiple LLMs under different prompting strategies. Figure 4 and Table 3 present the results. We highlight several key findings from the evaluation:
Split Model Method Subject Area Certification
miF maF miF maF
Public Filter LoRA 80.72 71.92 65.87 63.45
Qwen3-4B 0-shot 55.43 51.61 45.77 30.49
Qwen3-4B 4-shot 56.33 54.42 45.46 30.28
Qwen3-4B CoT 59.72 55.66 47.80 35.77
Private Filter LoRA 79.06 70.48 65.54 63.50
Qwen3-4B 0-shot 42.93 31.73 44.12 25.01
Qwen3-4B 4-shot 45.76 34.08 44.04 29.41
Qwen3-4B CoT 46.22 35.49 47.70 31.92
Table 4: Expertise Classification Performance
@@
Table 5: End-to-end RAG Performance and Ablation Study on ExpertRAG and Expert-CoT.
@@
Few-shot prompting improves accuracy up to a point. We varied the number of in‑context exemplars from 0 to 64 (See Appendix A.7.2) and observed that incorporating a small number of examples yields substantial gain over the zero‑shot baseline. However, adding examples beyond a certain point leads to diminishing or no improvement.
@@
Baseline LLMs underperform on easier questions. As shown in Figure 4a, across all models, we observe that performance is lowest for the EMR certification level (the most basic tier in the NREMT exam) and highest for the Paramedic level. This may be due to smaller data size for EMR level and the procedural nature of EMR questions, whereas Paramedic questions aligning more closely with the medical content seen during LLM pretraining.
@@
LLMs falter on the core NREMT domains. Figure 4b shows that models reliably answer “pharmacology” and “anatomy” questions but struggle in “pediatrics” and core areas such as “trauma”,“airway”, “operations”, and “cardiology”. One possible reason is subject area complexity. Questions in the former areas often need single-hop, fact-based queries solvable in a zero-shot manner, whereas the latter demand multi-hop reasoning and richer EMS knowledge.
Expertise Classification Performance
The performance of our Filter vs. LLM baselines (0-shot, 4-shot, and CoT) for expertise classification are shown in Table 4. Since subject area classification is a multi-label task and certification classification is a multi-class task, we report micro f1-score (miF) and macro f1-score (maF). Results show our Filter trained with LoRA with two classification heads significantly outperforms the baseline LLMs.
Expert-CoT Evaluation
To assess how domain expertise, injected via Expert-CoT, influences reasoning (RQ2), we compare Expert-CoT with different prompting strategies under multiple LLMs. As shown in Table 3, Expert-CoT help guide LLM reasoning. Integrating domain expert knowledge via CoT‑Expert guides reasoning towards appropriate context and consistently boosts CoT prompting performance by up to 2.05% across models. Also, using the predicted expertise attributes (Filter) vs. the ground-truth attribute annotations (GT) yields comparable performance for Expert-CoT, demonstrating the Filter’s strong performance, as also shown in Table 4.
Ablation Study on Expert-CoT and ExpertRAG
@@
We further evaluate the effect of injecting expertise into CoT and RAG (RQ2) using an ablation study with six configurations, as shown in Table 5: (1) No-RAG, (2) RAG+CoT with a standard global retriever on EMS PR and KB, (3) RAG+Expert-CoT, (4) ExpertRAG-GT+CoT, (5) ExpertRAG-GT+Expert-CoT, and (6) ExpertRAG-Filter+Expert-CoT. An ablation study on certification and subject area is presented in Appendix A.8. The best configuration outperforms the baseline by 4.59 / 3.69 points in Acc / F1 (See error analysis in Appendix A.9).
@@
Effect of PR and KB. (2) vs. (1) isolates the gain of adding PR and KB as retrieval documents with a standard global retriever. Results show KB brings more improvement than PR, and combing both yields the best performance.
@@
Effect of Expert-CoT. (3) vs. (2) (and (5) vs. (4)) ablates the additional gain from Expert-CoT on RAGs, showing that expertise-aware reasoning is better than standard reasoning.
@@
Effect of Expert-RAG. (4) vs. (2) (and (5) vs. (3)) ablate the impact of our Expert-RAG (FTR/RTF) compared to standard RAG with a global retriever. With ground-truth subject area and certification level, ExpertRAG consistently outperforms SOTA RAG baselines, highlighting the value of expertise-guided retrievers. Both FTR and RTF outperform global retrieval, with RTF achieving better performance.